# Cost-driven prunings for iterative solving of constrained routing problem with SRLG-disjoint protection

### v1

@@ I Introduction
Combination of the two problems gives the SRLG-disjoint DCR problem. To the best of our knowledge, the only existing algorithm to address directly this problem is CoSE-Pulse+, published in [22]. Note that although they call the problem SRLG-disjoint DRCR, Delay-Range Constrained Routing, their formulation does not consider the lower bound on delay and thus represents only the trivial case of delay-range. The work [22] pointed out, that solution of SRLG-disjoint DCR problem may require, as an intermediate step, to solve a single path DRCR problem with lower bound on delay. For this problem it proposed the Pulse+ algorithm, a modification of the Pulse algorithm, to handle the lower bound delay constraint. This work also published a dataset containing both real-world and randomly generated networks, for performance evaluation of algorithms solving both DRCR and the SRLG-disjoint DCR problems.
@@
We benchmark our algorithms on the original and extended dataset, providing a thorough comparison with Pulse+ and CoSE-Pulse+ algorithms and highlighting their advantages and disadvantages (Section IV). From the benchmark results and analysis, we conclude that our algorithms are better suited to the case when one needs to find as many feasible solutions as possible, meeting strict time requirements. On the other hand, CoSE-Pulse+ has the advantage in proving the non-existence of feasible solutions and better worst-case performance in certain scenarios.
@@ II-B SRLG-disjoint constrained routing problem
Here, $AP$ and $PP$ stands for Active Path and Protection Path. In a typical scenario, all traffic is routed by AP, while PP is used only in the case of AP failure. So, it is reasonable to minimize cost only for AP, letting PP be any available path.
The problem with a nontrivial lower bound on delay may arise in the following scenario. Different internet services have different requirements for connection quality [23]. There are tasks mentioned in the Introduction section that rely on limited and deterministic delay values. Other applications, such as distributed cloud office, can tolerate tens of milliseconds in delay without quality degradation, while others, like e-mail or file transfer, have pretty much no requirements for signal delay. When all these services occupy the same infrastructure, it may be reasonable to set an artificial lower bound on delay for non-critical applications to reserve low-latency connections for more valuable tasks.
@@
All SRLG-disjoint routing problems can be classified into non-trap and trapped instances, the latter being more challenging to solve. The notion of trap originates from a solution technique called Active Path First, or APF [24], [25], [26]. In this technique, the DRCR problem for AP is first solved, then all edges sharing SRLG with the found AP are removed from the graph, and delay bounds are updated according to $d_{diff}$. The DRCR problem is then solved again in the graph with removed edges for PP. If it succeeds, the solution satisfies all the requirements and is optimal. But in some cases the shortest candidate for AP does not have feasible disjoint pair, that could serve as PP. This is the so-called trap in the literature, and is of particular interest for research, since in this case simple APF approach is not enough. Traps can be avoidable and unavoidable, depending on whether or not a feasible pair of paths exists in principle. For the case of avoidable trap, some sophisticated strategy is required to find the solution [19]. For unavoidable traps, occasions arise when one needs to prove the non-existence of solutions.
@@ II-C Pulse algorithm
The successful implementation of Pulse algorithm has 3 main components: search order, pruning strategies and the action on the terminal node. Some attempts to improve search order were made in works [15], [16], [22] to further increase performance of the original Pulse for the CSP problem. [22] also proposed a change in pruning strategies to handle lower bounds on delay, producing the Pulse+ version of the algorithm. It differs from the original Pulse in two points: firstly, the dominance pruning is not applicable in this case, and secondly, tracking of visited nodes is required for each pulse to avoid cycles in the found path.
One weakness of Pulse / Pulse+ algorithm originates from the fact that optimality pruning comes into play only after the first feasible path was found. So, there is no pruning by cost for the search of the first feasible path. As shown in the next section, this first-found feasible solution is often far from optimality due to statistical reasons, because the majority of paths are situated at higher values of cost. In section III-B we address this problem, proposing to introduce iteratively updated heuristic upper-bounds on cost. This enables pruning by cost from the start of the search, thus effectively restricting the search space.
In section III-D we further develop this idea and provide an efficient technique for the solution of SRLG-disjoint DRCR problem, that uses modifications of Pulse+ as a subroutine. For that, we will modify the third component of the algorithm, namely the action at a terminal node. Thus we obtain three versions of the algorithm: the original Pulse+ searching for the shortest single constrained path, search for all AP candidates in given cost range and search for a feasible PP without optimality requirement. The pseudocode and discussion of all used versions of Pulse+ can be found in Appendix A.
@@ III Algorithms
As mentioned in the previous section, only two pruning strategies are applicable in the presence of a delay constraint: infeasibility and optimality pruning. While the first is deterministic, the second depends on how the search is carried out and the cost of the first found feasible path. Our innovation aims to use pruning by cost in a more controllable way. We can gain insight into this by analyzing the structure of search space with respect to the path cost.
@@ III-A Path distribution for DRCR problem
To better illustrate the intuition behind our algorithms, we describe the distribution of paths for a typical problem instance (Figure 1). This example is qualitatively representative for many problem cases. We provide more examples for other network topologies in the Appendix D. On the figure the path distribution relative to their cost for one particular task is shown. The data for this diagram is obtained in two steps. First, ordinary Pulse+ algorithm without optimality pruning is adopted to get the distribution of feasible paths. It takes about 40 minutes to finish. To count all paths including infeasible we use a modification of Pulse+ algorithm that is designed for all path search in cost corridors (see Appendix A-B) with relaxed delay constraint and termination when path count reaches $10^{8}$. It takes few hours to finish.
@@
The diagram shows that paths satisfying the $d_{up}$ constraint follow a bell-shaped distribution around some pretty high cost value. The total number of paths in this category is more than 4 billion, which is, in effect, the size of the initial search space. At the beginning, $current\_min\_cost$ is set to infinity, so the search is conducted among all paths that satisfy $d_{up}$ constraint (Figure 2 (a)). Since the DFS-like search in the original algorithm is unguided (basically, random), the most probable path to be found tends to lie in the region where feasible paths are most frequent. After that, the optimality pruning comes into play, installing the upper bound on cost and discarding the right-hand part of path distribution. However, due to the bell-like shape of the distribution, more frequent thus more probable to find paths still lie towards the higher values of cost. So, each next path to be found will most probably have cost value close to the current value of $current\_min\_cost$. Consequently, the value of $current\_min\_cost$ will be decreased only a little on each iteration, and the search process will resemble slow motion ”down the hill” by the feasible path distribution.
Figure 2: Dynamics of $current\_min\_cost$ in the Pulse algorithm. a) The distribution of feasible paths (blue) and search space that is left after infeasibility pruning (orange) in the process of Pulse algorithm updates. b) Actual cost dynamics for one typical task. c) Our proposal to estimate initial value of $current\_min\_cost$ based on the cost for the shortest unconstrained path
@@ III-B Bottom-top exploration approach to DRCR problem
For the single-path CSP problem, our work is a natural extension of the Pulse+ algorithm, which we call Bottom-Top Bounds Update (BTBU). Recall that the initialization step of the Pulse / Pulse+ algorithm calculates the shortest path tree from the target node to each node in the graph. So, before the Pulse algorithm starts, we have a reference lower-bound value for cost, which is the cost of the shortest path from the source to the target. The approach of Pulse+ is to set $current\_min\_cost=infinity$ and run the algorithm. Instead, we propose to place a judiciously chosen upper bound on the cost and search for feasible paths under this upper bound. If no such paths are found, we assume that the chosen upper bound is too tight and lies below the cost of the shortest feasible path. In this situation, we increase the upper bound according to a specific strategy and perform the search again until a feasible path is found. At first glance, it may seem ineffective because we may need many runs before finding the feasible solution. Nevertheless, all unsuccessful runs will be performed near the tail of the path distribution between the cost of the shortest unconstrained path and the cost of the shortest feasible path, meaning that very little time is spent exploring nearly empty search space. The initial estimate of the upper bound and the update rule for it must be chosen so that the upper bound at each iteration is not too loose to effectively restrict the search space, while at the same time keeping number of necessary algorithm reruns reasonably low before the area of feasible paths is reached.
@@ 7      $current\_min\_cost\ +=\ cost\_step$
We propose two approaches that were found to be highly effective in our experiments (Algorithms 1, 2). For the first one, the initial value of current_min_cost is two times larger than the cost of the shortest unconstrained path. On each iteration, if the search was unsuccessful, we increase the cost by another factor of two. For the second approach, we introduce a variable called $cost\_step$ which is given by some function of the cost of a graph edge e.g. the cost of the shortest edge, or average cost of an edge in the graph, or some other variations. We start from value of $current\_min\_cost$ just one single step higher than the cost of the shortest unconstrained path. When an iteration is unsuccessful, we multiply the step by 2, add to the $current\_min\_cost$ and run the Pulse+ search again.
In both algorithms we multiply the step by 2 on each iteration, thus increasing the bound in geometric progression. Such update rule was found the most effective in our experiments. Algorithm 1 is more minimalistic, since both the initial estimate for the upper bound and the update rule are based solely on the cost of the shortest unconstrained path. Algorithm 2 allows us to choose proper function $f(G)$, that may better match the particular network. In our experiments we used the cost of the shortest edge in the graph for the $f(G)$, or double that value.
@@ III-C Path distribution for SRLG-disjoint DCR problem
In this section we describe the intuition behind our BTCS algorithm for the SRLG-disjoint D(R)CR problem. Figure 3 shows the distribution of active paths for a typical instance of the problem. The network, used to generate this data is smaller than the one chosen for section III-A (only 197 nodes and 490 links), so, it is possible to plot the distribution of all paths from source to target for this task. We put the partial distributions for several other instances into Appendix D to show, that this picture in its essential features is indeed typical for the problem under consideration. Although the distribution of paths on Figure 3 may look similar to that on Figure 1, note that the legend of two figures is different.
Recall the problem statement for SRLG-disjoint DCR problem from section II-B. We are interested not only in the minimum cost of the active path (AP) but also in one that satisfies delay constraints and has a protection path (PP). Let us look at the problem from the perspective of AP while considering the existence of a disjoint PP as a property of each AP. We can divide all paths that connect the given source node to the target node into three categories. The first category contains infeasible paths that do not satisfy delay constraints. The second category is paths that satisfy constraints but do not have a disjoint PP. The third category is paths that satisfy constraints and have at least one disjoint PP. The shortest path in the last category and the corresponding disjoint PP are the paths we are looking for. We call such a pair of shortest feasible AP together with disjoint PP the solution of the problem.
At the left end of the distribution we can notice, that the shortest (lowest cost) candidate for the active path does not have disjoint pairs. This is exactly the previously discussed trap situation. Contrary to some rather anecdotal examples presented in related literature (see, for example, [22] and discussion in Appendix D), we notice that the shortest path having protection pair is also located close to the shortest one, near the left tail of the distribution. Therefore the natural idea is again to restrict the search in the area of low costs, instead of searching across the whole space of feasible paths.
@@ III-D Solving SRLG-disjoint DRCR problem: corridor search
Here we describe our proposed algorithm, that uses several adaptations of Pulse+ as intermediate steps. We call it BTCS-Pulse+, bottom-top corridor search. BTCS-Pulse+ is exact in the sense that it guarantees optimality of the found solution. In terms of performance, it has both advantages and disadvantages compared to COSE-Pulse+ (please see discussion in Section IV). The main features of the technique are depicted on Figure 4 and the pseudocode is presented in Algorithm 3.
@@
In order to find the solution, we first run a single iteration of Pulse+. It efficiently handles the delay constraints and returns the shortest feasible AP candidate. It cannot account for the existence of a disjoint PP for this candidate, which we have to check separately. For that, it removes from the graph all edges that share an SRLG with edges in the found AP and update the delay lower and upper bounds according to the delay difference constraint based on the actual delay of the found AP. Removal of the links that share SRLG with AP guarantees that AP and PP will not share any SRLGs. And update of the upper and lower bounds of delay is necessary to ensure that the difference in delay between AP and PP satisfies the delay difference constraint. After this, we run Pulse+ again on the modified graph and with new delay range constraint to find a PP. If PP is successfully found, then this pair of AP and PP is indeed the solution of the problem.
@@
If this first AP candidate does not have a PP, we shall start searching for other candidates. Since we are interested in finding the shortest protected AP, it is natural to start from the already found one which has the lowest cost and go up to the paths of higher costs, searching for feasible AP candidates and checking them one by one for the existence of disjoint PP, until such candidate is found. This movement from the lowest cost path candidates to the higher costs is why we call our algorithm bottom-top search.
@@
We take the cost of the already found AP candidate as a starting point and divide the whole range of costs above it into a sequence of reasonably small corridors. The width of the corridor is a hyperparameter of the algorithm. It can be estimated based on the costs of the edges in the graph or tuned by hand. We will discuss the choice of corridor width in more detail later. The algorithm will explore those corridors one by one, starting from the one with the lowest cost and moving in the bottom-top direction.
@@
In the main loop of the algorithm we explore one particular cost corridor, that is given by current lower and upper bounds on cost. To efficiently find all feasible AP candidates in a corridor, we propose a modification of the Pulse+ algorithm, that uses the upper side of the corridor for pruning the search space above and stores all paths within the corridor instead of searching for the shortest one. We can see the details of this Pulse+ implementation in Appendix A-B. This version of the Pulse+ algorithm uses cost and delay in a similar way, since delay range is given in the problem statement, and cost range corresponds to the corridor. So, we use the upper bound of the cost corridor for pruning by cost together with the original infeasibility pruning by the upper bound on delay. And instead of updating the current minimal cost, as it is done in the original Pulse+ to find the optimal path, we save all paths, that satisfy upper and lower bounds on delay and cost. Thanks to the pruning by both delay and cost, the effective search space is kept to be sufficiently small (it corresponds to the left tail of the feasible distribution on Figure 3). If the corridor width is chosen to be not too wide, the number of AP candidates found in each corridor is also reasonably low.
@@
Having found all the AP candidates within current cost corridor, we sort them by increasing value of cost, take them one by one and search for a disjoint PP for each of them in a similar procedure to what was described above for the first found AP candidate: removing from the graph all edges, that share SRLG with current AP candidate, updating upper and lower bounds on delay based on the delay of current AP and delay difference constraint, then using Pulse+ to search for PP. If a disjoint PP is found for one of the AP candidates, then this pair of paths is returned as the solution for the problem, and algorithm stops. If there is no disjoint PP for all AP candidates in the current corridor, then it proceeds to the next corridor by updating the lower and upper bounds on cost and search for AP candidates again.
@@
This procedure is repeated until the solution is found. Since in a typical case the AP, that has a disjoint pair, lies close to the lower end of the distribution (See Fig. 3 and Appendix D), this search typically requires only a few corridors that contain only a few feasible AP candidates, which makes the search fast and efficient. Take for example the task presented on Fig 3. If the width of the cost corridor is chosen to be 10, then each corridor corresponds to a single bin of the diagram. In this case, 11 corridors will be explored, before the shortest AP that has a disjoint PP is found (the one shown in blue in the corridor between cost values 975 and 985). Five of those corridors are empty, and total number of AP candidates necessary to check (the yellow ones having costs lower, than the first blue) will be slightly above 10.
@@
2calculate $starting\_cost$ and first AP candidate by Pulse+ algorithm /* See Section II-C and Appendix A-A */
@@
7      $min\_cost$ = cost of the shortest edge in the graph
8      $cost\_step=f(G)$ /* Select the corridor width based on the costs of the edges in the graph or choose it as a hyperparameter */

### v2

@@ I Introduction
Combination of the two problems gives the SRLG-disjoint DCR problem. To the best of our knowledge, the only existing algorithm to address directly this problem is CoSE-Pulse+, published in [22]. The work [22] pointed out, that solution of SRLG-disjoint DCR problem may require, as an intermediate step, to solve a single path DRCR (Delay-Range Constrained Routing) problem with both lower and upper bounds on delay. For this problem it proposed the Pulse+ algorithm, a modification of the Pulse algorithm, to handle the lower bound delay constraint.
@@
We benchmark our algorithms on the wide range of networks of different type and size, providing a thorough comparison with Pulse+ and CoSE-Pulse+ algorithms and highlighting their advantages and disadvantages (Section IV). From the benchmark results and analysis, we conclude that our algorithms are better suited to the case when one needs to find as many feasible solutions as possible, meeting strict time requirements. On the other hand, CoSE-Pulse+ has the advantage in proving the non-existence of feasible solutions and better worst-case performance in certain scenarios.
@@ II-B SRLG-disjoint constrained routing problem
Here, $AP$ and $PP$ stands for Active Path and Protection Path. In a typical scenario, all traffic is routed by AP, while PP is used only in the case of AP failure. So, it is reasonable to minimize cost only for AP, letting PP be any available path. We note that this is not the only possible choice for the objective. Another important example is so called min-sum problem, that requires to minimize total cost of the two disjoint paths. This class of problems is typically approached via modifications of Suurballe’s heuristic [23], [18]. Our problem thus classifies as min-min problem.
The problem with a nontrivial lower bound on delay may appear in the following scenario. Different internet services have different requirements for connection quality [24]. There are tasks mentioned in the Introduction that rely on limited and deterministic delay values. Other applications, such as distributed cloud office, can tolerate tens of milliseconds in delay without quality degradation, while others, like e-mail or file transfer, have pretty much no requirements for signal delay. When all these services occupy the same infrastructure, it may be reasonable to set an artificial lower bound on delay for non-critical applications to reserve low-latency connections for more valuable tasks.
@@
All SRLG-disjoint routing problems can be classified into non-trap and trapped instances, the latter being more challenging to solve. The notion of trap originates from a solution technique called Active Path First, or APF [25], [26], [27]. In this technique, the DRCR problem for AP is first solved, then all edges sharing SRLG with the found AP are removed from the graph, and delay bounds are updated according to $d_{diff}$. The DRCR problem is then solved again in the graph with removed edges to find PP. If it succeeds, the solution satisfies all the requirements and is optimal. But in some cases the shortest candidate for AP does not have feasible disjoint pair, that could serve as PP. This is the so-called trap in the literature, and is of particular interest for research, since in this case simple APF approach is not enough. Traps can be avoidable and unavoidable, depending on whether or not a feasible pair of paths exists in principle. For the case of avoidable trap, some sophisticated strategy is required to find the solution [19]. For unavoidable traps, sometimes proof of infeasibility is needed.
@@
We also note that in case of unconstrained problem and planar network topology polynomial-time algorithms are already known for certain types of SRLG distribution, see for example [28], [29]. In current work we consider constrained problems, and do not leverage any particular network topology. On the contrary, in Section IV we test our algorithms on as much diverse dataset as we can, to highlight all aspects of its performance.
@@ II-C Pulse algorithm
The successful implementation of Pulse algorithm has 3 main components: search order, pruning strategies and the action on the terminal node. Some attempts to improve search order were made in works [15], [16], [22] to further increase performance of the original Pulse for the CSP problem. [22] also proposed a change in pruning strategies to handle lower bounds on delay, producing the Pulse+ version of the algorithm. It differs from the original Pulse in two points: firstly, the dominance pruning is not applicable in this case, and secondly, tracking of visited nodes is required for each pulse to avoid cycles in the found path. The pseudocode of Pulse+ algorithm is presented in Algorithms 1 and 2.
1
input : Graph G, source $s$, target $t$, delay range $[d_{low},d_{up}]$.
2
[…]
5Initialization(G, $t$). /* calculate min cost and min delay trees from each node to target with the help of Dijkstra or A* algorithm */
[…]
7Pulse_plus($P^{*}$, $P$, $s$, $visited\_nodes$, $current\_cost$, $current\_delay$, $current\_min\_cost$)
8return current_min_cost, $P^{*}$
Algorithm 1 Pulse+ algorithm
1
output : void
2
3if $n==t$ and $d_{low}\leq d\leq d_{up}$ and $c<current\_min\_cost$ then
4       $P^{*}=P$
5      $current\_min\_cost=c$
6      return
7 end if
8
9for $(nk)$ in egress links of node $n$ do
10       if $k$ in $visited\_nodes$ then continue;
11       if $d+d(nk)+d(P^{min\_delay}_{k\rightarrow t})>d_{up}$ then continue;
13
15 end for
Algorithm 2 Pulse_plus() recursive function
One weakness of Pulse / Pulse+ algorithm originates from the fact that optimality pruning comes into play only after the first feasible path was found. So, there is no pruning by cost for the search of the first feasible path. As shown in the next section, this first-found feasible solution is often far from optimality due to statistical reasons, because the majority of paths are situated at higher values of cost. In section III-B we address this problem by introducing iteratively updated heuristic upper-bounds on cost. This enables pruning by cost from the start of the search, thus effectively restricting the search space.
In section III-D we further develop this idea and provide an efficient technique for the solution of SRLG-disjoint DRCR problem, that uses modifications of Pulse+ as a subroutine. For that, we will modify the third component of the algorithm, namely the action at a terminal node. Thus we obtain three versions of the algorithm: the original Pulse+ searching for the shortest single constrained path, also an algorithm to search for all AP candidates in given cost range and another one to search for a feasible PP without optimality requirement. The pseudocode and discussion of these versions can be found in Appendix A.
@@ III Algorithms
As mentioned in the previous section, only two pruning strategies are applicable in the presence of a delay constraint: infeasibility and optimality pruning. While the first is deterministic, the second depends on the cost of the first found feasible path and how fast it is found. Our innovation aims to use pruning by cost in a more controllable way. We can gain insight into this by analyzing the structure of search space with respect to the path cost.
@@ III-A Path distribution for DRCR problem
To illustrate the intuition behind our algorithms, we describe the distribution of paths for a typical problem instance (Figure 1). This example is qualitatively representative for many problem cases. We provide more examples for other network topologies in the Appendix D. On the figure the path distribution relative to their cost for one particular task is shown. The data for this diagram is obtained in two steps. First, ordinary Pulse+ algorithm without optimality pruning is adopted to get the distribution of feasible paths. It takes about 40 minutes to finish. To count all paths including infeasible we use a modification of Pulse+ algorithm that is designed for all path search in cost corridors (see Appendix A-B) with relaxed delay constraint and termination when path count reaches $10^{8}$. It takes few hours to complete.
@@
The diagram shows that paths satisfying the $d_{up}$ constraint follow a bell-shaped distribution around some pretty high cost value. The total number of paths in this category is more than 4 billion, which is, in effect, the size of the initial search space. At the beginning, $current\_min\_cost$ is set to infinity, so the search is conducted among all paths that satisfy $d_{up}$ constraint (Figure 2 (a)). Since the DFS-like search in the original algorithm is unguided (basically, random), the most probable path to be found tends to lie in the region where feasible paths are most frequent. After that, the optimality pruning comes into play, installing the upper bound on cost and discarding the right-hand part of path distribution. However, due to the bell-like shape of the distribution, more frequent thus more probable to find paths still incline towards the higher values of cost. So, each next path to be found will most probably have cost value close to the current value of $current\_min\_cost$. Consequently, the value of $current\_min\_cost$ will be decreased only a little on each iteration, and the search process will resemble slow motion ”down the hill” by the feasible path distribution.
Figure 2: Dynamics of $current\_min\_cost$ in the Pulse algorithm. a) The distribution of feasible paths (blue) and search space that is left after infeasibility pruning (orange) in the process of Pulse algorithm updates. b) Actual cost dynamics for one typical task. c) Our proposal to estimate initial value of $current\_min\_cost$ based on the cost of the shortest unconstrained path
@@ III-B Bottom-top exploration approach to DRCR problem
For the single-path CSP problem, our work is a natural extension of the Pulse+ algorithm, which we call Bottom-Top Bounds Update (BTBU). Recall that the initialization step of the Pulse / Pulse+ algorithm calculates the shortest path tree from the target node to each node in the graph. So, before the Pulse algorithm starts, we have a reference lower-bound value for cost, which is the cost of the shortest path from the source to the target. The approach of Pulse+ is to set $current\_min\_cost=infinity$ and run the algorithm. Instead, we propose to place a judiciously chosen upper bound on the cost and search for feasible paths under this upper bound. If no such paths are found, the chosen upper bound is too tight and lies below the cost of the shortest feasible path. In this case, the upper bound is updated according to a specific strategy and the search is repeated again until a feasible path is found. At first glance, it may seem ineffective because many re-runs may be needed before finding the feasible solution. Nevertheless, all unsuccessful runs will be performed near the tail of the path distribution between the cost of the shortest unconstrained path and the cost of the shortest feasible path, meaning that very little time is spent exploring nearly empty search space. The initial estimate of the upper bound and the update rule for it must be chosen so that the upper bound at each iteration is not too loose to effectively restrict the search space, while at the same time keeping number of necessary algorithm reruns reasonably low before the feasible region is reached.
@@ 7      $current\_min\_cost\ +=\ cost\_step$
Two approaches were found to be effective in our experiments (Algorithms 3, 4). For the first one, the initial value of current_min_cost is two times larger than the cost of the shortest unconstrained path. On each iteration, if the search was unsuccessful, the cost is again multiplied by 2. For the second approach, we introduce a variable called $cost\_step$ which is given by some function of the cost of a graph edge e.g. the cost of the shortest edge, or average cost of an edge in the graph, or some other estimation. The search starts from value of $current\_min\_cost$ just one single step higher than the cost of the shortest unconstrained path. When an iteration is unsuccessful, the step is multiplied by 2, added to the $current\_min\_cost$ and the Pulse+ search runs again.
In both algorithms the step is multiplied by 2 on each iteration, thus the bound increases in geometric progression. Such update rule was found the most effective in our experiments. Algorithm 3 is more minimalistic, since both the initial estimate for the upper bound and the update rule are based solely on the cost of the shortest unconstrained path. Algorithm 4 allows to choose proper function $f(G)$, that may better match the particular network. In our experiments we used the cost of the shortest edge in the graph for the $f(G)$, or double that value.
@@ III-C Path distribution for SRLG-disjoint DCR problem
In this section we describe the intuition behind our BTCS algorithm for the SRLG-disjoint D(R)CR problem. Figure 3 shows the distribution of active paths for a typical instance of the problem. The network, used to generate this data is smaller than the one chosen for section III-A (only 197 nodes and 490 links), so it is possible to plot the distribution of all paths from source to target for this task. We put the partial distributions for several other instances into Appendix D to show, that this picture in its essential features is indeed typical for the problem under consideration. Although the distribution of paths on Figure 3 may look similar to that on Figure 1, note that the legend of two figures is different.
Recall the problem statement for SRLG-disjoint DCR problem from section II-B. We are interested not only in the minimum cost of the active path (AP) but also in one that satisfies delay constraints and has a protection path (PP). Let us look at the problem from the perspective of AP while considering the existence of a disjoint PP as a property of each AP. We can divide all paths that connect the given source node to the target node into three categories. The first category contains infeasible paths that do not satisfy delay constraints. The second category has paths that satisfy constraints but do not have a disjoint PP. The third category has paths that satisfy constraints and have at least one disjoint PP. The shortest path in the last category and the corresponding disjoint PP are the paths we are looking for. We call such a pair of shortest feasible AP together with disjoint PP the solution of the problem.
At the left end of the distribution we can notice, that the shortest (lowest cost) candidate for the active path does not have disjoint protection. This is the previously discussed trap situation. Contrary to some rather exaggerated examples presented in literature (see, for example, [22] and discussion in Appendix D), we notice that the shortest path having protection pair is also located close to the shortest one, near the left tail of the distribution. Therefore the natural idea is again to restrict the search in the area of low costs, instead of searching across the whole space of feasible paths.
@@ III-D Solving SRLG-disjoint DRCR problem: corridor search
Here we describe our proposed algorithm, that uses several adaptations of Pulse+ as intermediate steps. We call it BTCS-Pulse+, bottom-top corridor search. BTCS-Pulse+ is exact in the sense that it guarantees optimality of the found solution. In terms of performance, it has both advantages and disadvantages compared to COSE-Pulse+ (see discussion in Section IV). The main features of the technique are depicted on Figure 4 and the pseudocode is presented in Algorithm 5.
@@
The initial stage of the algorithm consists of a single run of Pulse+. It efficiently handles the delay constraints and returns the shortest feasible AP candidate. It cannot account for the existence of a disjoint PP for this candidate, which has to be tested separately. For that, the algorithm removes from the graph all edges that share an SRLG with the found AP and updates the delay lower and upper bounds according to the delay difference constraint based on the delay of the found AP. Removal of the links that share SRLG with AP guarantees that AP and PP will not share any SRLGs. And update of the upper and lower bounds of delay is necessary to ensure that the difference in delay between AP and PP satisfies the delay difference constraint. After this, another run of Pulse+ is performed on the modified graph with new delay range constraint to find a PP. If PP is successfully found, then this pair of AP and PP is the solution of the problem, and algorithm stops.
@@
If this first AP candidate does not have a PP, the algorithm starts to search for other candidates. Since we are interested in finding the shortest protected AP, it is natural to start from the already found one which has the lowest cost and gradually move to the paths of higher costs, searching for feasible AP candidates and checking them one by one for the existence of disjoint PP, until such candidate is found.
@@
The cost of the initially found shortest AP candidate serves as a starting point. The whole range of costs above it is divided into a sequence of reasonably small corridors. The corridor width is a hyperparameter of the algorithm. It can be estimated based on the costs of the graph edges, tuned by hand or pretrained on historic data for the same network. The algorithm will explore those corridors one by one, starting from the smallest cost and moving in the bottom-top direction.
@@
The main loop of the algorithm explores one particular cost corridor, that is given by current lower and upper bounds on cost. To efficiently find all feasible AP candidates in a corridor, we propose a modification of the Pulse+ algorithm, that uses the upper side of the corridor to prune the search space above, and enumerates all feasible paths within the corridor instead of searching for the shortest one. We put the details of this Pulse+ implementation in Appendix A-B. This version of the Pulse+ algorithm uses cost and delay in a similar way, as delay range is given in the problem statement, and cost range corresponds to the corridor. The upper bound of the cost corridor serves for pruning by cost together with the original infeasibility pruning by the upper bound on delay. And instead of updating the cost bound, as it is done in the original Pulse+ to find the optimal path, this version store all paths, that satisfy upper and lower bounds on delay and cost. Thanks to the pruning by both delay and cost, the effective search space is kept to be sufficiently small (it corresponds to the left tail of the feasible distribution on Figure 3). If the corridor width is chosen to be small enough, the number of AP candidates found in each corridor is also reasonably low.
@@
The AP candidates found in current cost corridor are sorted by cost in ascending order and taken one by one to search for a disjoint PP for each of them. It is done in a similar manner to what was described above for the first found AP candidate: removing from the graph all edges, that share SRLG with current AP candidate, updating upper and lower bounds on delay based on the delay of current AP and delay difference constraint, then using Pulse+ to search for PP. If a disjoint PP is found for one of the AP candidates, then this pair of paths is the solution for the problem, and algorithm stops. If there is no disjoint PP for all AP candidates in the current corridor, then the algorithm proceeds to the next corridor by updating the lower and upper bounds on cost and looks for AP candidates again.
@@
This procedure is repeated until the solution is found. Since in a typical case the AP, that has a disjoint pair, is situated close to the lower end of the distribution (See Fig. 3, Appendix D, and also Fig. 6 (d), (e) and corresponding figures in Appendix E), this search typically requires only a few corridors that contain only a few feasible AP candidates, which makes the search fast and efficient. Take for example the task on Fig 3. If the width of the cost corridor is equal to $10$, then each corridor corresponds to a single bin of the diagram. In this case, $11$ corridors will be explored, before the shortest AP that has a disjoint PP is found (the one shown in blue in the corridor between cost values $975$ and $985$). Five of those corridors are empty, and total number of AP candidates necessary to check (the yellow ones having costs lower, than the first blue) will be slightly above $10$.
@@
2calculate $starting\_cost$ and first AP candidate by Pulse+ algorithm /* See Section II-C */
@@
7      $cost\_step=f(G)$ /* Compute the corridor width based on the costs of the edges in the graph or choose it as a hyperparameter */